### v1

# VLA-Pruner: Temporal-Aware Dual-Level Visual Token Pruning for Efficient Vision-Language-Action Inference

Abstract
Vision-Language-Action (VLA) models have shown great promise for embodied AI, yet the heavy computational cost of processing continuous visual streams severely limits their real-time deployment. Token pruning—keeping salient visual tokens and dropping redundant ones—has emerged as an effective approach for accelerating Vision-Language Models (VLMs), offering a solution for efficient VLA. However, these VLM-specific token pruning methods select tokens based solely on semantic salience metrics (e.g., prefill attention), while overlooking the VLA’s intrinsic dual-system nature of high-level semantic understanding and low-level action execution. Consequently, these methods bias token retention toward semantic cues, discard critical information for action generation, and significantly degrade VLA performance. To bridge this gap, we propose VLA-Pruner, a versatile plug-and-play VLA-specific token prune method that aligns with the dual-system nature of VLA models and exploits the temporal continuity in robot manipulation. Specifically, VLA-Pruner adopts a dual-level importance criterion for visual token retention: vision-language prefill attention for semantic-level relevance and action decode attention, estimated via temporal smoothing, for action-level importance. Based on this criterion, VLA-Pruner proposes a novel dual-level token selection strategy that adaptively preserves a compact, informative set of visual tokens for both semantic understanding and action execution under given compute budget. Experiments show that VLA-Pruner achieves state-of-the-art performance across multiple VLA architectures and diverse robotic tasks.
1 Introduction
Recent advances in robot learning have led to remarkable progress in developing generalist policies capable of acting across diverse tasks and environments [35, 13, 62, 7, 63, 22, 4, 25, 9]. Vision–Language–Action (VLA) models [7, 63, 22, 23, 4, 25, 30], which benefit from the multimodal understanding and transfer ability of Vision–Language Models (VLMs) [33, 3], have emerged as a promising pathway toward general-purpose robot policy. Leveraging large-scale real-world robotic datasets [40], VLA models can perceive visual scenes, understand natural language instructions, and directly execute low-level actions to complete complex tasks across diverse scenarios. However, in real-time deployment, VLA models must process continuous visual streams within dynamic environments and thus struggle with heavy computational overhead. To mitigate this, existing works have explored acceleration techniques such as model lightweighting [51], quantization [41], and early-exit [59] for VLA. Despite effectiveness to some extent, these methods require either architectural modifications or retraining, therefore hindered by poor generalizability.
Previous research [8, 47, 2] has shown that visual representations (i.e., visual tokens) in Vision-Language Models (VLMs) are highly redundant. Consequently, visual token pruning, which retains the most salient tokens while pruning the rest, has become a general and effective approach to improve VLM inference efficiency [2, 8, 60, 47, 57, 53, 64] and shows promise for general VLA acceleration. These methods typically define importance criteria for tokens, such as attention scores [8, 60, 64], gradient information [37, 36] or feature diversity [2] to quantify the significance of visual tokens, and less important tokens are pruned at inference time. By reducing visual redundancy, they yield substantial compute savings with minimal accuracy loss on standard VLM tasks such as image captioning [1, 44], VQA [20, 34], and video understanding [54, 28].
Despite recent progress, existing token pruning methods typically rely solely on semantic salience metrics measured during context prefilling—e.g., prefill attention distributions in FastV [8] or text-to-vision cross-attention scores in SparseVLM [60]—to rank visual tokens and prune the lowest-scoring ones. While such semantic-only criteria can be effective for VLMs, they are misaligned with the intrinsic dual-system nature of VLA models [18, 19, 48, 11]. Specifically, VLA inference couples high-level semantic understanding, task planning with low-level action execution, which entails distinct visual requirements between its vision-language prefill and action decode stages; correspondingly, we observe clearly different attention patterns across these two stages (see Sec. 3.2). Consequently, current methods tend to bias token retention towards semantic cues, thereby discarding critical details for action generation and significantly degrading VLA performance, especially at high pruning ratios (see Fig. 1), as also observed in prior work [55]. Thus, existing VLM-specific token pruning strategies are fundamentally ill-suited for VLA models, highlighting the need for a VLA-specific method.
To bridge this gap, we propose VLA-Pruner, a versatile plug-and-play VLA token pruning approach that aligns with the dual-system nature of VLA models and exploits temporal continuity in robot manipulation. VLA-Pruner offers two principal insights for VLA token pruning. First, VLA-Pruner adopts a dual-level importance criterion for visual token retention: vision-language prefill attention scores to quantify semantic-level relevance (following prior works [8, 57, 37, 60]), and action decode attention scores to quantify action-level importance. A critical challenge to achieve this is that decode attention is unavailable at prefill time. Fortunately, we observe that temporal continuity of robot manipulation [55] manifests in VLA action attention patterns evolving: top-attended visual patches overlap substantially across consecutive timesteps (see Sec. 3.2). Leveraging this continuity, we practically estimate the current action-to-vision attention from recent decode attentions via statistical temporal smoothing techniques (e.g., Exponential Moving Average [15, 14]). Second, to ensure VLA performance, our novel dual-level token selection strategy avoids fragile attention score fusion that incurs redundancy and bias. Instead, drawing on the minimal-redundancy-maximal-relevance (mRMR) principle [42], we adopt a patch-wise scheme: we first take the union of patches salient at either semantic or action level to maximize relevance, and then filter them by reducing redundancy, yielding a compact yet informative visual token set that meets the given compute budget. This innovative design reduces computational overhead and adaptively preserves essential information for both semantic understanding and action execution. VLA-Pruner is training-free, serving as a plug-and-play acceleration module for most state-of-the-art VLA architectures [22, 24, 4]. We evaluate VLA-Pruner on robotic manipulation across two simulated environments (LIBERO [32] and SIMPLER [27]) and three VLA models (OpenVLA [22], OpenVLA-OFT [24], and [4]). VLA-Pruner delivers substantial efficiency gains (up to 1.8 speedup) with minor drops in task success rate. Even at 87.5% prune ratio, VLA-Pruner can maintain reasonable performance. Moreover, thanks to precise redundancy removal, VLA-Pruner can improve model performance at the 50% prune ratio. Finally, we demonstrate its real-world applicability by deploying it on a 6-DoF xArm6 robot, achieving practical speedups in real-time scenarios.
2 Related Work
2.1 Vision-Language-Action Models (VLA)
Large-scale vision–language models (VLMs) have propelled multimodal learning by incorporating visual perception and language reasoning [26]. Building on these advances, Vision–Language–Action (VLA) models [22, 4, 24, 7, 63] introduce an action modality, enabling end-to-end motor control. These models typically adopt large VLM backbones [49] and are fine-tuned on robotic data [40]. For action generation, implementations range from discretizing actions into language-like tokens (e.g., OpenVLA [22]) to attaching diffusion-policy heads (e.g., [4]), typically relying on action-to-vision cross-attention for fine-grained interaction. Despite their effectiveness on robot manipulation tasks such as object retrieval and assembly [32, 27], VLA models remain computationally intensive, hindering real-time deployment on resource-constrained platforms.
2.2 Visual Token Pruning for VLMs
Visual token pruning [8, 60, 2, 47, 57] is a widely adopted strategy to reduce visual redundancy in VLMs. One line of methods defines semantic importance criteria to rank and prune visual tokens, e.g., FastV [8] (using early-layer attention), SparseVLM [60] (using text–to-vision cross-attention), DivPrune [2] (using diversity-driven selection), and HoloV [64] (using crop-wise attention and diversity). Another line of calibration-based approaches (e.g., FitPrune [57], VTW [31]) select pruning layers/ratios by analyzing model outputs on a calibration set. However, the first line relies mainly on semantic salience and overlooks the action-dependent requirements of VLA, leading to performance drops (see Sec. 3.2). The second line depends on offline calibration that is unfriendly to resource-constrained VLA deployment and still hinges on semantic salience. These limitations motivate a VLA-specific pruning method.
2.3 Training-Free Acceleration for VLA Models
Visual token pruning has recently been integrated into training-free VLA acceleration frameworks [58]. For example, EfficientVLA [58], SP-VLA [29], and SpecPrune-VLA [50] perform structured acceleration that incorporates token pruning. However, these methods still use semantic salience to retain tokens. Thus, they rely on additional modules (e.g., diffusion-feature cache [56], lightweight generation module [29], action-aware controller [50], or layer-reduction [56, 50]) to balance performance and speed, limiting their generalizability and effectiveness. A recent method, VLA-Cache [55], exploits temporal continuity by caching static visual token features, opening a promising direction. However, its cache-based mechanism is less efficient than direct pruning and still depends on text-to-vision cross-attention to preserve task-relevant tokens in a coarse-grained manner. The above limitations stem from overlooking the dual-system nature of VLA models (see Sec. 3.2). In this paper, we show that token pruning can achieve significant VLA acceleration without sacrificing performance.
3 Preliminary Analysis
3.1 Background
VLA Inference.
Vision-Language-Action (VLA) models process sequential inputs of the form into across timesteps . Here, represents the input visual observations, denotes language task instructions, and denotes the predicted action tokens, which can be detokenized into executable robotic actions. The text input is tokenized into text tokens . The vision encoder processes into image features, which are projected into visual tokens using a projector. These tokens, along with a proprioceptive token representing robot state, form the Transformer input. Since self-attention scales quadratically with token count, the (generally ) visual tokens dominate inference cost.
VLA Attention Breakdown.
Like VLMs, VLA models rely on the attention mechanism of transformer for token interactions. Without loss of generality, we describe the single-head attention below. At each layer , the model takes previous-layer hidden states of visual tokens
and textual tokens
as inputs. For mathematical simplicity, we omit time/layer superscripts and treat the proprioceptive token as part of the vision-language context. These inputs are transformed into queries , keys , and values using linear projections. Then, the shared vision-language context for attention and interaction is formed as: . We break down VLA inference into two stages:
(1) Vision-Language prefill Stage.
In this stage, queries come from the text and vision tokens:
The attention matrix is computed by
Attention score for each input visual patch is computed as
(2) Action decode Stage. In this stage, queries are action tokens ( for autoregressive [22], for chunk-based [4]). Attention is computed as
For the autoregressive case, attention vectors from decoding steps are concatenated into the final attention matrix. The action-to-vision attention score in this stage is
and quantifies each patch’s significance to semantic understanding [55, 60] and action execution, respectively.
3.2 Motivating Observation for VLA-Pruner
We systematically delve into the intrinsic characteristics of VLA inference (Fig. 2). Our findings identify the fundamental limitations of current VLM-specific token pruning methods for VLA models, which motivate our VLA-Pruner.
Motivation 1: Visual token prefilling dominates computational overhead for VLA models.
As noted in previous works [60, 8], incorporation of visual tokens inevitably introduces substantial memory and computational overhead when compared to LLMs. For VLA models, the number of processed visual tokens (typically 256 per frame, where is the number of camera views) is often an order of magnitude larger than that of text tokens (30–50) and action tokens (7–56), making input visual stream prefilling the primary contributor to the computational cost. Since visual representations are highly redundant [2, 8, 60], visual token pruning shows great promise for VLA acceleration.
Motivation 2: Visual token pruning must account for the dual-system nature of VLA models.
The efficacy of VLA models stems from their inherent dual-system ability to integrate high-level semantic understanding and task planning with low-level action execution [13, 18, 19, 48, 11]. Our key insight is that this inherent duality manifests directly in VLA model inference, resulting in distinct visual requirements for its two stages: vision-language prefill and action decode. To investigate this, we analyzed the attention distributions of OpenVLA [22] on the LIBERO benchmark [32] (Fig. 2). The results reveal that VLA inference demonstrates distinct attention patterns: prefill attention exhibits a broad, semantic distribution, while action decode attention shows localized focus essential for precise motor control, as exemplified in Fig. LABEL:fig:prefill_attn and Fig. LABEL:fig:decode_attn. Quantitatively, the overlap ratio of top-attended visual tokens between vision–language prefilling and action decoding averages 50% and often falls below 30% (Fig. 2). Attention scores are averaged across all Transformer layers. Existing pruning strategies overlook this crucial attention divergence, thereby risking the premature removal of tokens vital for action generation but less salient during prefilling, particularly at high compression ratios. This oversight is critical, since removing even a few high-attention tokens is known to severely degrade performance [61, 16]. We also provide a qualitative illustration in Fig. 7.
Motivation 3: VLA token pruning should exploit temporal continuity in robotic manipulation.
Since vision token prefilling dominates computational overhead (Motivation 1), token pruning must operate during early-layer prefilling [8]. However, while an ideal strategy should consider both semantic and action-level token importance (Motivation 2), action decode attention scores (practical proxy for action-level importance) are unavailable at prefill time. Fortunately, VLA robot manipulation exhibits temporal continuity [55]. We observe that action-to-vision attention maps show strong short-term temporal correlation: the top-attended visual tokens during action decoding at timesteps and show consistently high overlap ratios (Fig. 2 show ratios averaged across LIBERO tasks and a representative rollout). This temporal continuity makes statistical temporal smoothing techniques [52, 14, 21] an practical and effective proxy for estimating current action-to-vision attention.
4 Methodology
4.1 Problem Formulation
We define the token pruning problem as follows: given a set of visual tokens with and a target subset size , the goal is to select a subset that preserves information critical for VLA inference. Formally, we define a mapping that selects an index subset with , and constructs the retained visual tokens as . The objective is to identify a mapping function that minimizes the difference in the model’s output before and after pruning:
where and . denotes the conditional generation probability of VLA model. denotes a loss function that measures the difference in the model’s output with and without pruning, and indicates the number of retained tokens. We conceptually decompose into vision-language prefill and action decode . Let denote all hidden representations of textual and visual tokens after prefilling. Informally,
where also drops the corresponding indices from the hidden visual representations . As noted in Motivation 2, optimizing with a single-level objective, i.e., semantic-level or action-level , leads to suboptimal pruning. To address this, we propose VLA-Pruner.
4.2 The Proposed VLA-Pruner
VLA-Pruner adopts a (a) dual-level token importance criterion that incorporates both semantic-level and action-level importance for visual token retention, and proposes a corresponding (b) dual-level token selection strategy to adaptively process dual-level relevance and redundancy for token retention. The overall pipeline is illustrated in Fig. 3.
(a) Dual-level Token Importance Criterion
For VLMs, previous works [8, 60] have demonstrated that prefill or text-to-vision attention scores can effectively quantify significance of vision tokens. In VLA case, as discussed before (Sec. 3.2), both semantic-level and action-level visual requirements should be satisfied. To this end, VLA-Pruner adopts a dual-level token importance criterion: (1) prefill attention scores (Eq. 2) to quantify semantic-level relevance and (2) action decode attention scores (Eq. 3) to quantify action-level importance. As discussed in Motivation 3, we build on the temporal continuity of VLA manipulation to estimate current action attention scores over image patches. This formulates a time-series forecasting task [5], where is estimated from historical trend . Classical training-free forecasting methods include averaging, autoregressive models, and exponential smoothing [6, 5, 52]. Among them, the Exponential Moving Average (EMA) [21, 15, 14] provides a simple yet effective way to estimate temporal trends. By emphasizing recent observations while discounting older ones, EMA reduces noise, captures temporal trends, and has been widely applied in domains such as finance, economics, meteorology, and robotic manipulation [17, 43, 10, 39]. Directly applying EMA to our task takes this form:
where is the smoothing factor, denotes the smoothed prediction at the previous timestep, and represents the corresponding observation result.
Decaying Window Average Mechanism. For VLA manipulation, where short-term context is more critical [55], we instead propose a simpler and more intuitive decaying window average mechanism. This method considers recent timesteps and applies an explicit exponential decay. Formally, we smooth action attention scores from previous steps, , and compute the estimated score for current step as their weighted sum:
where defines the window size, and controls the decay rate. This formulation provides a simple yet effective mechanism to estimate current action attention scores. For brevity, we omit time superscript in following discussion.
Remark 1.
Temporal smoothing methods can reliably estimate action attention dynamic trends but may fail under abrupt changes, such as target switching. For instance, in the LIBERO-Long task “put both the alphabet soup and the tomato sauce in the basket”, action attention from putting the soup does not ground the tomato sauce well. Potential remedies include lightweight predictors for attention shifts or mechanisms to bypass pruning when large prediction–observation discrepancies occur. To ensure generalizability and flexibility, we instead tackle this fundamental challenge through our dual-level token selection strategy.
(b) Dual-Level Token Selection Strategy.
Based on the dual-level token importance criterion, an intuitive approach to retain tokens is: normalizing semantic and action attention scores ( and ), taking a weighted sum, and selecting the top- tokens by the weighted score. However, this scheme suffers from fundamental limitations. First, it introduces a sensitive weighting hyperparameter that must be carefully tuned, compromising robustness and generalizability. Second, it ignores redundancy both within and across stages. On one hand, as noted in prior work [2], attention-based pruning often retains similar tokens. On the other hand, score-level fusion tends to retain tokens that appear salient in both prefilling and decoding, yet not essential to either stage. For example, in a robotic arm, intermediate arm segments may appear salient across both stages, whereas the end-effector (critical for execution) can be undervalued. Finally, it implicitly assumes that the action attention estimate is always reliable, which may not hold under abrupt attention shifts (see Remark 1).
Patch-wise Combine-then-Filter. To avoid these limitations, we draw on the minimal-Redundancy-Maximal-Relevance (mRMR) principle [42], an information-theoretic paradigm for feature selection. mRMR proposes to combine both Max-Relevance and Min-Redundancy principles, reducing the redundant features while keeping the relevant features for the task. Inspired by this, our dual-level token selection strategy follows a patch-wise combine-then-filter mechanism. We organize the procedure into three phases:
(1) Dual-level Top-k Selection.
We begin by constructing two candidate token subsets via dual-level top-k selection:
where is the target token budget. This phase identifies salient tokens (indices) for high-level semantic understanding () and low-level action execution (), respectively.
(2) Relevance-Maximization Pooling. We then combine these two subsets to form a candidate pool:
This subset constructs a candidate pool that contains tokens essential for dual-level information requirements, maximizing two-stage task relevance. Due to attention divergence (Motivation 2), is generally larger than .
(3) Redundancy-Minimization Filtering. This phase aims to reduce feature redundancy within the candidates to satisfy the target budget . We achieve this by maximizing the diversity of the final selected subset. Following prior work [2], we formulate it as a Max–Min Diversity Problem (MMDP) [45] within candidate pool , where the goal is to find a subset of size that maximizes the minimum pairwise distance between its elements:
where is any subset of with elements, and are input image patch embeddings indexed by any . The distance is defined as the cosine distance:
A solution for the MMDP problem 8 is a subset of that minimizes redundancy between elements. MMDP can be solved by several exact and heuristic methods [46, 38]. To facilitate efficiency, we use a greedy algorithm that iteratively add tokens whose minimum distance to the retained set is maximal, until reaching elements. To escape local optima, we initialize the first token with maximal second-nearest distance. We detail the step (3) and our overall token selection method in Algorithm 1 and 2 (see Appendix A).
Remark 2.
Our method is a heuristic solution for mRMR objective: the first two steps maximize task relevance, while the final step minimizes feature redundancy. The challenge in Remark 1 is addressed by grounding abrupt shifts by prefill attention and retaining them via diversity maximization.
Remark 3.
Step (3) shares similarities with DivPrune [2], which reduces redundancy through diversity-based pruning. However, DivPrune does not incorporate relevance maximization step, thus fail to identify pivotal tokens. By combining both Max-relevance and Min-redundancy, we retain token subsets that are both diverse and better matched to dual-level information requirements within VLA inference.
(c) Implementation Details.
VLA-Pruner prunes visual tokens at Transformer layer K within VLA model (K is set to 3 [8]). We select tokens using last-layer prefill attention and temporal smoothing of action-to-vision attention averaged over the latter half layers (to reduce noise) from past timesteps. Pruned tokens are dropped from the rest prefill layers. To adapt to flash-attention [12], adaptation techniques in SparseVLM [60] can be adopted. For decaying window average, window size and decay rate are empirically set as 3 and 0.8. We warm-start VLA-Pruner with steps to record sufficient action attention history. VLA-Pruner is a plug-and-play acceleration module for VLA architectures with action-to-vision cross-attention, a mechanism used by state-of-the-art VLA models [22, 24, 4], including (i) autoregressive policies [22, 7, 63, 63] (e.g., OpenVLA [22]), (ii) action-chunk decoders (OpenVLA-OFT [24]), and (iii) diffusion-head policies such as [4], for which we average action attention scores over its flow-matching steps. We provide analysis of computational complexity in Appendix A.2.
[…]
5 Experiments
We evaluate VLA-Pruner in both simulation and real-world settings. In simulation setting, we evaluate VLA-Pruner on open-source VLA models: OpenVLA [22], OpenVLA-OFT [24] and [4], using the LIBERO benchmark [32] and SIMPLER environment [27]. All experiments are conducted on an NVIDIA RTX 4090 GPU.
5.1 Experimental Setup
Baselines. We compare VLA-Pruner with a comprehensive suite of training-free accelerators under given compute budgets. Baselines include state-of-the-art token pruning methods (FastV [8], SparseVLM [60], and DivPrune [2]) and a VLA-specific method VLA-Cache [55]. We provide more details of baselines and VLA models in Appendix B.1. EfficientVLA [56] adopts token prune strategies similar to prior work (using attention and diversity), we include results of it and more token pruning baselines in Appendix LABEL:app:more_baselines. Evaluation Protocol. All methods are evaluated under identical given pruning budgets. To ensure a comprehensive comparison, our experiments include diverse pruning ratios (50%, 75%, 87.5%). For VLA-Cache, we set the same token reuse ratio. Our evaluation metrics mainly include: task success rate (%), inference latency (ms) and FLOPs(T).
5.2 Evaluation Benchmarks
In this section, we introduce the evaluation benchmarks, with more details presented in Appendix B.2.
LIBERO.
The LIBERO benchmark [32] comprises four suites (Spatial, Object, Goal, and Long) that probe complementary aspects of manipulation generalization. Each suite contains 10 distinct tasks with 50 evaluation episodes per task (500 total episodes per suite). Following OpenVLA [22] and OpenVLA-OFT [24], we adopt their standard evaluation setup (action chunks as 8 for OpenVLA-OFT) and official weights to ensure comparability. To verify cross-architecture generality, we also include results of diffusion-head based model [4].(see Appendix LABEL:app:pi0_libero)
SIMPLER.
The SIMPLER simulators [27] provide two complementary settings aimed at closing sim-to-real gaps: visual matching (VM) and variant aggregation (VA). We report evaluation results on three tasks and their corresponding settings: Move Near, Pick Coke Can (under VA), and Open/Close Drawer (under VM). We choose these task–setting pairs where the vanilla OpenVLA attains non-trivial success rates, ensuring a reasonable comparison.
Real Robot Evaluation.
To assess VLA-Pruner’s performance in real-world scenarios, we conduct experiments using a 6-DoF xArm6 robotic arm equipped with a parallel gripper. We run four manipulation tasks involving diverse objects (100 trials for each task): (1) Can Stack: grasp the can and stack it on top of the box; (2) Cup Pour: pick up the cup and pour the small balls inside the cup into the box; (3) Cube Place: grasp the cube and place it into the box; (4) Bottle Place: grasp the bottle and place it into the box.
5.3 Main Results
Results on LIBERO.
We evaluate VLA-Pruner on OpenVLA (fine-tuned) and OpenVLA-OFT for LIBERO benchmark. Table 4.2 reports success rate on the four suites under token-retention ratios of 50%, 25%, and 12.5%. We also report a relative accuracy metric that measures pruning-induced error, defined as the performance preservation ratio (vanilla model as 100%). VLA-Pruner achieves the best success rate and relative accuracy across all settings among training-free accelerators. Even at a 12.5% retention ratio, it maintains 88.9% and 88.27% relative accuracy on OpenVLA and OpenVLA-OFT, respectively, outperforming baselines by up to 34.39%. At a 50% retention ratio, VLA-Pruner can even improve success rate, thanks to precise noise filtering that stabilizes policy execution. This is evident on LIBERO-Long, where it achieves 4.2% success rate improvement. DivPrune benefits VLM inference but degrades VLA performance since robot manipulation needs more localized visual details. VLA-Cache achieves competitive results due to temporal awareness but fails at high compression ratios because its text-to-vision attention selection strategy is coarse-grained. Fig. 4 visualizes more details of VLA-Pruner performance varying under different pruning ratios. We present results on in Appendix LABEL:app:pi0_libero.
Method Move Near Pick Coke Can Open/Close Drawer Overall OpenVLA 54.0% 52.8% 49.5% 52.1% (100%) (100%) (100%) (100%) FastV 38.7% 41.9% 33.7% 38.1% (71.7%) (79.4%) (68.1%) (73.1%) VLA-Cache 41.2% 40.6% 38.8% 40.2% (76.3%) (76.9%) (78.4%) (77.2%) VLA-Pruner 52.4% 50.1% 48.8% 50.4% (97.0%) (94.9%) (98.6%) (96.8%)
Results on SIMPLER.
We evaluate cross-environment generalization on the SIMPLER environment. Table 2 compares VLA-Pruner with acceleration baselines on OpenVLA at 75% compression ratio. VLA-Pruner best preserves accuracy, showing robust transfer across environments.
Results on Real Robot.
We evaluate real-world performance on a physical 6-DoF xArm6 arm controlled by OpenVLA-OFT. Fig. 5 reports results under 75% token-pruning ratio. VLA-Pruner attains the highest relative accuracy, underscoring its practicality for on-robot deployment.
Efficiency Analysis.
We report inference efficiency metrics in Table 4.2, including inference latency ( or ) and FLOPs(T). At same pruning ratios, VLA-Pruner achieves comparable speedups to visual token–pruning methods. Its latency is 5 ms higher than FastV (due to distance computations in redundancy filtering), yet it achieves significantly better performance. At the 25% retention ratio, VLA-Pruner outperforms token pruning baselines at 50% retention, while being 16% faster and using 20% fewer FLOPs. VLA-Pruner is faster and more compute-efficient than VLA-Cache. VLA-Pruner requires storing historical attention, but it introduces a negligible increase in Max GPU memory (see Appendix LABEL:app:more_results:memory).
5.4 More Results
Ablation Studies.
To validate our (a) dual-level importance criterion and (b) patch-wise dual-level token selection strategy, we ablate three variants: (1) prefill-only, which selects tokens by prefill attention; (2) action-only, which selects tokens by action attention; and (3) score-fusion, which ranks tokens by the weighted sum of prefill and action attention scores, with a carefully fine-tuned weight. As shown in Fig. 6, prefill-only preserves high-level semantics but hinders fine-grained low-level control, whereas action-only improves short-horizon control but sacrifices semantic grounding and task planning, and thus neither maintains VLA performance well. Score-fusion remains inferior to the prefill-only setting. We also ablate the temporal decay window size in Fig. 6, which demonstrates the robustness of our method and reveals the short-term temporal continuity of VLA manipulation. When , it reduces to using the last timestep’s action attention to guide pruning, leading to performance degradation. We further ablate varying decay rate and pruning layer in the Appendix LABEL:app:more_results:ablations.
Qualitative Visualization.
We provide a qualitative case study of VLA-Pruner in Fig. 7. VLA-Pruner preserves essential localized visual information for precise action execution, while maintaining broad semantic understanding.
6 Conclusion
We propose VLA-Pruner, a training-free token pruning method for efficient VLA inference. To align with the VLA models’ dual-system nature of high-level semantic understanding and low-level action execution, VLA-Pruner employs a dual-level token importance criterion that integrates semantic-level relevance and action-level importance. Based on this criterion, our dual-level token selection strategy adaptively retains visual tokens. Across diverse robotic manipulation tasks and multiple VLA architectures, VLA-Pruner reduces computation while preserving performance. We discuss our limitations in Appendix C.
Supplementary Material
Appendix A Method Details
A.1 Algorithm Outlines
We detail VLA-Pruner in Algorithm 1 (min-redundancy filtering) and Algorithm 2 (overall dual-level token selection).
A.2 Complexity Analysis
Following FastV [8], SparseVLM [60] and DivPrune [2], we estimate the computational requirement of VLA-Pruner in terms of Transformer FLOPs, and focus on how pruning visual tokens changes the effective sequence length seen by each layer. As in prior work [2], we count the FLOPs of multi-head attention and feed-forward networks (FFNs) in the backbone (the overhead of token selection as negligible compared to the main model FLOPs.) Let be the total number of Transformer layers in the VLA backbone and the index of the pruning layer. Denote the numbers of text and visual tokens before pruning by and (), and let be the full sequence length. After VLA-Pruner keeps a fraction of the visual tokens, the sequence length becomes . Following [8], the FLOPs of attention and FFN for one layer with sequence length , hidden size , and FFN intermediate size can be approximated as
Without pruning, the total FLOPs are
With VLA-Pruner, the first layers still use the full sequence , while the remaining layers use the shortened sequence , giving
Thus, the theoretical FLOPs ratio is
Since the attention term in is quadratic in sequence length, the reduction scales roughly with in layers after . This matches the empirical FLOPs ratios in Tab. 4.2: VLA-Pruner reduces FLOPs to about , , and of the original model at , , and token retention, respectively. Temporal smoothing over a short window of action-to-vision attention has complexity per timestep, and the greedy Max–Min filtering over the dual-level candidate pool of size is per prefill. In practice, these components contribute little to the total wall-clock time, as these operations are invoked only once during the prefill stage. Finally, Eq. (13) shows that VLA-Pruner offers a flexible way to trade accuracy for efficiency: the retained-token ratio . Smaller reduces FLOPs quadratically in layers after , yielding higher speedup, while larger recovers the original model’s performance. In our experiments, we report results at , but VLA-Pruner can be used at any intermediate ratio to meet a given compute or latency target.
Appendix B Experiments Details
B.1 Setting up Details
B.1.1 VLA Model Details
We consider three representative VLA architectures in our experiments: OpenVLA, OpenVLA-OFT, and . All models take RGB observations and a natural language instruction as input, and produce robot actions end-to-end. OpenVLA. OpenVLA [22] is a 7B-parameter open-source vision–language–action model built by fine-tuning a Prismatic VLM (Llama 2 language backbone with DINOv2 and SigLIP visual encoders) on robot manipulation episodes from the Open X-Embodiment dataset. The model represents actions as discrete tokens and decodes them autoregressively, enabling a single policy to control multiple embodiments out-of-the-box. OpenVLA serves as our base autoregressive VLA, and we follow its official checkpoints and evaluation settings on LIBERO. OpenVLA-OFT. OpenVLA-OFT [24] is an “Optimized Fine-Tuning” variant of OpenVLA that improves both control performance and inference efficiency. It augments the original decoder with parallel decoding and action chunking, and switches to a continuous action representation trained with a simple L1 regression objective. This substantially boosts success rates on LIBERO while increasing action-generation throughput. To be noted, OpenVLA-OFT utilizes bidirectional attention, meaning that during inference, there is already an action-to-vision cross-attention. However, we observe that during the early-layer prefill phase, the attention still focuses on semantic understanding, making action-to-vision less significant. In deeper layers, the action-to-vision attention starts to exhibit more distinct properties. This allows our method to effectively enhance token pruning performance by utilizing past timestep action-to-vision cross-attention at the later layers to guide current token pruning. Model. [4] is a prototype generalist robot policy that augments a pretrained vision–language model (with a PaliGemma backbone) with a continuous-action flow-matching head. Built on top of a large VLM to inherit Internet-scale semantic knowledge, is trained on diverse real-robot data and is designed for high-frequency control in dexterous manipulation tasks. The model decodes continuous motor commands via a flow-matching process with 10 integration steps. We use as a diffusion-based VLA backbone to evaluate the generality of VLA-Pruner. To capture the most informative action-to-vision interactions, we record the action-to-vision cross-attention averaged over the latter half of these integration steps, where the cross-attention becomes more stable and reliable.
B.1.2 Acceleration Baseline Details
We briefly summarize the baseline acceleration methods considered in our experiments.
FastV. FastV [8](ECCV24) accelerates large vision–language models by pruning redundant visual tokens based on early-layer prefill attention. Its core observation is that many image patches receive consistently low attention in deeper layers while contributing substantially to computation. FastV defines an attention-based importance score during the prefill stage and drops a fixed fraction of low-score visual tokens after a chosen layer. It is training-free and purely semantic, but does not account for the action-dependent requirements of VLA models.
SparseVLM. SparseVLM [60](ICML25) also targets visual token redundancy in VLMs, but uses text-to-vision cross-attention as the importance signal. Tokens that receive low cross-attention from text queries are considered less relevant and pruned. Compared to FastV, SparseVLM focuses directly on aligning visual tokens with language semantics, yet still operates only at the semantic level and ignores the distinct visual needs of action decoding in VLA.
DivPrune. DivPrune [2](CVPR25) addresses the tendency of attention-based methods to select similar and redundant tokens. Instead of ranking tokens by attention scores, it formulates token selection as a Max–Min Diversity Problem and chooses a subset of visual tokens that are maximally dissimilar in feature space, thereby reducing redundancy while maintaining coverage. However, DivPrune is agnostic to the dual-level relevance of visual tokens in VLA prefilling and action decoding, and its diversity objective fall short for VLA tasks where localized details are more critical.
VLA-Cache. VLA-Cache [55](Neurips25) is a training-free acceleration method tailored for VLA models. It exploits temporal continuity in robotic manipulation by identifying visually static tokens whose features change little across timesteps, and reuses their computation via a KV-cache. At the same time, it uses text-to-vision cross-attention to keep task-relevant tokens fully computed. This cache-based mechanism reduces repeated computation over static background regions, but its token selection remains coarse-grained, which does not explicitly consider action-to-vision attention as we do. To ensure a fair comparison, we set the dynamic token limit based on similarity to be ( is the token budget), and use text-to-vision score to exclude of the most relevant tokens. These settings follow the default setup of VLA-Cache codebase.
EfficientVLA. EfficientVLA [56](Neurips25) proposes a structured acceleration framework for VLA models that combines visual token pruning, layer reduction, and diffusion temporal feature caching. Its token-pruning component leverages attention and diversity heuristics inherited from VLM pruning, while additional modules (i.e., layer skipping and diffusion caching) are used to achieve comparable acceleration. In contrast, VLA-Pruner is a single, plug-and-play token pruning module. These approaches are therefore complementary: EfficientVLA and other VLA acceleration methods can be adopted orthogonally with VLA-Pruner, while VLA-Pruner achieves strongest token-prune performance.
| Method | Spatial | Object | Goal | Long | Avg. | FLOPs | Speedup |
| VLA-Pruner | 85.4 | 82.5 | 78.4 | 51.8 | 74.5 | 39.80% | 1.633 |
| EfficientVLA (w/o layer) | 85.1 | 81.6 | 76.5 | 51.7 | 73.7 | 58.12% | 1.334 |
| EfficientVLA | 84.9 | 80.3 | 76.1 | 50.3 | 72.9 | 41.30% | 1.491 |
| VTW | 74.7 | 76.2 | 72.3 | 48.9 | 68.0 | 55.17% | 1.325 |
| FitPrune | 82.4 | 80.1 | 74.8 | 50.3 | 71.9 | 47.71% | 1.397 |
B.2 Benchmark Details
Fig. 7(a) shows the evaluation simulation environments and Fig. 7(b) presents our real-world xArm6 robot setup.
LIBERO.
LIBERO [32] is a Robosuite-based benchmark for lifelong robot manipulation that groups 130 language-conditioned tasks into four suites: Spatial, Object, Goal, and Long. LIBERO-Spatial varies only the spatial configuration of otherwise identical objects (e.g., pick up the black bowl on the stove and place it on the plate). LIBERO-Object varies the manipulated object while keeping the scene layout and goal template fixed (e.g., pick up the bbq sauce and place it in the basket). LIBERO-Goal keeps both the objects and spatial layout fixed but changes the goal description (e.g., open the middle drawer of the cabinet). LIBERO-Long corresponds to the long-horizon subset (10 tasks) of LIBERO-100, where each episode consists of multi-stage instructions such as opening a drawer, inserting an object, and then rearranging other items(e.g., put both the alphabet soup and the cream cheese box in the basket).
SIMPLER.
SIMPLER [27] is a simulation suite built to closely track the real Google robot environment while allowing controlled appearance variation. It provides two evaluation settings. The Visual Matching (VM) setting keeps backgrounds, lighting, distractors, and camera poses close to the real setup so that rendered observations visually resemble real robot scenes and better predict real-world performance. The Variant Aggregation (VA) setting starts from the same base scenes but procedurally perturbs nuisance factors such as background textures, illumination, number and placement of distractor objects, table material, and camera viewpoints, creating a broad distribution shift for testing robustness. For the Google robot configuration, SIMPLER defines four canonical manipulation tasks: “pick coke can”,”pick object”, “move near”, “open/close Drawer”, and “place in closed drawer”. In our experiments, we follow the official simulator settings and success metrics, and report results on three representative tasks—Pick Coke Can, Move Near, and Open/Close Drawer—under VM or VA setting.
Real robot.
As shown in Figure 7(b), our real-world system is built on a 6-DoF xArm6 manipulator equipped with two Intel RealSense D435i RGB cameras: one mounted on the wrist for close-range observations and the other fixed for capturing global scene context. The robot is controlled through the official xArmAPI. We collect real-world demonstrations following the LeRobot 2.1 data specification. Each episode consists of synchronized observations recorded at 30Hz, including two RGB image streams. The proprioceptive state is represented by the robot’s absolute joint angles, and actions are stored using the same absolute joint angles to ensure consistent replay and supervision. All demonstrations are gathered using this unified multi-view setup, which provides both global scene context and fine-grained manipulation details that are essential for learning robust visuomotor policies.
Pruning layer . We compare results of different pruning layer values in Table 6. VLA-Pruner details. We show ratio trends at different pruning ratios within a robotic task in Fig. 10.
| Method | TFLOP (ratio %) | Spatial | Object | Goal | Long | Avg |
| K=2 | 37.90 | 69.7 | 74.1 | 55.3 | 34.8 | 58.48 |
| K=3 (Ours) | 39.77 | 85.4 | 82.5 | 78.4 | 51.8 | 74.53 |
| K=4 | 41.98 | 85.5 | 82.3 | 78.1 | 52.0 | 74.48 |
| K=5 | 43.08 | 83.9 | 81.7 | 77.9 | 51.4 | 73.73 |
B.6 More Visualization
We provide detailed visualization of VLA-Pruner token prune results across LIBERO tasks under diverse ratios in Fig. 11. VLA-Pruner maintains both local details for action execution and global information for semantic understanding and task planning.
Appendix C Limitations and Future work.
Our main contribution is to point out the dual-system nature of VLA inference and to propose a dual-level importance criterion that jointly accounts for semantic and action-level relevance. However, our current estimation of action-level importance relies on a heuristic temporal smoothing scheme with a fixed window and decay rate. While this works well for most tabletop settings, its benefits may diminish in highly dynamic scenarios (e.g., egocentric wrist-camera views), where rapid viewpoint shifts and object motion can violate the short-term temporal continuity assumption and blur sharp changes in action-to-vision attention. A promising direction for future work is to replace the fixed temporal smoothing with an adaptive prediction module that adjusts its effective window based on motion cues or feature changes (e.g., magnitude of visual dynamics or action-conditioned attention variance), or to learn a lightweight temporal attention network that directly predicts per-token importance from recent history. Such adaptive mechanisms may further improve pruning quality in dynamic environments while retaining the efficiency advantages of VLA-Pruner.